You put $250 in the bank for 5 years at 12%.

If interest is added at the end of the year, how much will you have in the bank after one year?

Calculate the amount you will have in the bank at the end of year two and continue to calculate all the way to the end of the fifth year.

Show your Math
Answer #1
a) amount you will have in the bank at the end of year one
PV 250
r 0.12
n 5
FV= PV*(1+r)^n
Where,
FV= Future Value
PV = Present Value
r = Interest rate
n= periods in number
= $250*( 1+0.12)^5
=250*1.76234
= $440.59
b) amount you will have in the bank at the end of year two
FV= PV*(1+r)^n
= $250*( 1+0.12)^2
=250*1.2544
= $313.6
c) all the way to the end of the fifth year
FV= PV*(1+r)^n
= $250*( 1+0.12)^5
=250*1.76234
= $440.59